# PMGDA: A Preference-based Multiple Gradient Descent Algorithm

### 14 Feb 2024 v1

@@ Abstract
It is desirable in many multi-objective machine learning applications, such as multi-task learning and multi-objective reinforcement learning, to find a Pareto optimal solution that can exactly match a given preference of decision-makers. These problems are often large-scale with available gradient information but cannot be handled very well by the existing algorithms. To tackle this critical issue, this paper proposes a novel predict-and-correct framework for locating the exact Pareto optimal solutions required by a decision maker. In the proposed framework, a constraint function is introduced in the search progress to align the solution with a user-specific preference, which can be optimized simultaneously with multiple objective functions. Experimental results show that our proposed method can efficiently find exact Pareto optimal solutions for standard benchmarks, multi-task, and multi-objective reinforcement learning problems with more than thousands of decision variables.
@@ I Introduction
Recently, multi-objective optimization concepts and methodologies have received growing attention in the machine learning community since many machine learning applications can be naturally modeled as multi-objective optimization problems  [1, 2, 3]. For example, a face recognition system has to balance prediction accuracy on different attributes of the face [4, 5]. A recommendation system needs to balance several metrics, such as accuracy, novelty, and diversity of products [6, 7]. Very often, no single best solution can optimize all these objectives simultaneously. Instead of a single optimal solution, there exists a set of Pareto-optimal solutions with different optimal trade-off ratios.
Modern machine learning systems usually use deep neural networks as the core model. Multi-objective optimization problems in these machine learning systems [2, 8, 9, 5, 4] often share the following characteristics:
@@
TABLE I: Comparison with Other Methods
@@
 EPO [14] Exact
(1) Only for minimization tasks.
(2) Only for positive objectives.
(3) Precision is bad on various problems.
 COSMOS [4] Approximate Exact Solution quality dependent on hyperparameter $\mu$ and preferences.
 PMTL [8] Sector Areas Challenging to apply for $m>2$ objectives due to complex sector divisions.
 mTche [15] Exact
(1) Inefficient for problems with more than two objectives ($m>2$).
(2) Convergence is bad on complex reinforcement learning problems.
@@
Both Inequality and Equality Constraints.
Ideal for exact and ROI constraints.
Similar to EPO/PMTL, PMGDA needs to solve a linear/quadratic programming problem.
Due to these characteristics, gradient-based methods are particularly well-suited for solving multi-objective optimization problems in machine learning. It is, however, very costly to find a set of distinct Pareto optimal solutions (e.g., deep neural networks) to approximate the entire Pareto front accurately. Some efforts have been made to design multi-objective gradient algorithms to find one or a few Pareto solutions [2, 8, 16, 3, 17] with successful applications on multi-task learning, image classification, face recognition, and other scene understanding tasks [18, 2] based on the Multiple Gradient Descent Algorithm (MGDA) [19]. However, a significant limitation of most of these gradient algorithms [2, 8, 19, 20] is their lack of consideration for the decision maker’s (DM’s) preferences.
With a very limited number of solutions (only one or a few), it is of great importance to develop algorithms for finding Pareto optimal solutions satisfying user’s specialized demands. Some initial studies have explored approaches for finding Pareto objective solutions that align with user preferences. For instance, Pareto Multi-Task Learning (PMTL) [8] identifies Pareto solutions within specific sector regions, and the COSMOS method [4] incorporates a penalty term to approximate a Pareto solution near a preference vector. The Exact Pareto Optimization (EPO) method from Mahapatra et al. [14] focuses on locating solutions along a preference vector, claiming to find an ‘exact’ solution. However, these methods are still nascent and have their own disadvantages. For example, PMTL [8] struggles with efficiency for more than three objectives, and COSMOS [4] is limited to approximate solutions. EPO employs a complex and specific search mechanism to divide objectives into different sets with different treatments. This approach sometimes fails to find the preferred solutions  [^2] If the linear programming problem defined in EPO fails, EPO relies on optimizing a linear combination of objectives. and performs poorly with more than two objectives. A comprehensive comparison with previous methods is presented in Table I.
Different from previous works, this paper introduces a general Preference-based Multiple Gradient Descent Algorithm, termed PMGDA, to effectively address user-specific preferences within the MGDA framework [19]. The preference assignment in PMGDA is flexible and general, which includes but is not limited to the exact Pareto solution finding approach in EPO [14]. The proposed PMGDA has a predict-and-correct approach to optimize the preference and all objective functions simultaneously. This paper makes the following main contributions:
@@
    We propose a novel framework for identifying Pareto solutions that align with a wide range of general preference functions, where finding exact Pareto solutions is a special case. Our approach is the first to effectively find Pareto optimal solutions that meet specific region of interest (ROI) [21, 22] requirements.
@@
    We rigorously test PMGDA on well-known synthetic challenges, deep multi-task learning, and multiobjective reinforcement learning problems. PMGDA surpasses previous approaches in terms of solution precision and computational efficiency for finding exact Pareto optimal solutions, while it is the first approach to successfully find Pareto solutions that satisfy the ROI constraint.
@@ II-A Notations
In this paper, the following notations are used (refer to Table II): Vectors are denoted using bold letters (e.g. ${\bm{v}}$). Scalars are denoted using non-bold letters. A matrix is denoted by an uppercase letter (e.g., $G$). The index set $\{1,\ldots,m\}$ is denoted by $[m]$. We use three symbols to compare the relationship between two vectors ${\bm{x}}$ and ${\bm{y}}$:
For two vectors, ${\bm{x}}$ and ${\bm{y}}$, ${\bm{x}}\prec{\bm{y}}$ if ${\bm{x}}_{i}<{\bm{y}}_{i}$ for all $i\in[m]$. ${\bm{x}}\preceq{\bm{y}}$ if ${\bm{x}}_{i}\leq{\bm{y}}_{i}$ for all $i\in[m]$. ${\bm{x}}\preceq_{\text{strict}}{\bm{y}}$ if there exists $j\in[m]$ such that ${\bm{x}}_{j}<{\bm{y}}_{j}$ and ${\bm{x}}_{i}\leq{\bm{y}}_{i}$ for all $i\in[m]\setminus j$. In this article, ${\bm{x}}_{i}$ denotes the $i$-th entry of vector ${\bm{x}}$, and ${\bm{x}}^{(i)}$ denotes the $i$-th vector in a set of vectors.
TABLE II: The notation table
@@
$\Delta^{m-1}$ The (m-1)-simplex, $\left\{{\bm{y}}|\sum_{i\in[m]},{\bm{y}}_{i}=1,{\bm{y}}_{i}\geq 0\right\}$.
@@
$\hat{\bm{x}}$ $\hat{\bm{x}}=\frac{{\bm{x}}}{\left\|{\bm{x}}\right\|}$.
$\bm{\mu},\nu$ The Lagrange multipliers.
@@ II-B Basic MOO Definitions
A multi-objective optimization problem can be written as:
$\min_{{\bm{\theta}}\in\Theta\subset\mathbb{R}^{n}}{\bm{L}}({\bm{\theta}})=({\bm{L}}_{1}({\bm{\theta}}),{\bm{L}}_{2}({\bm{\theta}}),\ldots,{\bm{L}}_{m}({\bm{\theta}})),$ (1)
where ${\bm{\theta}}$ is a $n$-D decision variable, and $\Theta$ is the decision space. Very often, no single solution can optimize all the objectives at the same time. Instead, there exists a set of solutions, called Pareto optimal solutions, which can represent different optimal trade-offs among the objectives. The following MOO definitions are used in this paper. We have the following definition:
@@
A solution is Pareto optimal if no other solution belongs to $\Theta$ dominates it. The set of all the Pareto optimal solutions is called the Pareto set (PS); its image in the objective space is called the Pareto front ($\mathcal{T}$), $\mathcal{T}=f\circ\mathtt{PS}$.
@@
For objectives ${\bm{L}}_{1},\ldots,{\bm{L}}_{m}$ under the constraint $h({\bm{\theta}})$, a solution ${\bm{\theta}}$ is deemed Pareto stationary if it satisfies the following condition: there are no vectors ${\bm{\mu}}_{1},\ldots,{\bm{\mu}}_{m}$, and $\nu$ (not all zero) for which
@@
(a) Exact Pareto solution
(b) Pareto solutions with a region of interest (ROI)
Figure 1: Preference-based MOOs.
A solution is Pareto stationary means that, this solution can not be locally improved, which is a necessary condition for Pareto solution. Since a Pareto set could contain more than one single solution, and the possible infinite Pareto solutions are not comparable (e.g., non-dominated) with each other, therefore, it is necessary to make a further decision among the Pareto solutions [24]. Preference-based MOO aims to search for a Pareto optimal solution that satisfies a user-defined preference function $h({\bm{\theta}})=0$. A direct application is to search the so-called ‘exact’ Pareto solutions.
@@
Let ${\bm{\lambda}}=({\bm{\lambda}}_{1},\ldots,{\bm{\lambda}}_{m})$ be a positive preference vector, i.e., $\sum_{i=1}^{m}{\bm{\lambda}}_{i}=1$ and ${\bm{\lambda}}_{i}>0$, $\forall i\in[m]$. A solution ${\bm{\theta}}$ is called an ‘exact’ Pareto solution [14] for the preference ${\bm{\lambda}}$ if this solution is Pareto optimal and satisfies
$\frac{{\bm{L}}_{1}({\bm{\theta}})}{{\bm{\lambda}}_{1}}=\frac{{\bm{L}}_{2}({\bm{\theta}})}{{\bm{\lambda}}_{2}}=\ldots=\frac{{\bm{L}}_{m}({\bm{\theta}})}{{\bm{\lambda}}_{m}}.$ (3)
Another application of preference-based MOO is that, in many cases, particularly when dealing with large problem sizes, the user may only be interested in a specific region of interest (ROI) on the Pareto front [21]. This falls under the category of preference-based multi-objective optimization (MOO) [22]. Previous literature has defined the ROI using the reservation point $\overline{\bm{y}}$ and the aspiration point $\underline{\bm{y}}$. The ROI can be represented as $\mathtt{ROI}=\{\bm{y}|\underline{\bm{y}}\preceq\bm{y}\preceq\overline{\bm{y}}\}$ [25]. Illustrations of ‘exact’ solution and ROI are shown in Figure 1.
@@ II-C The Multiple Gradient Descent Algorithm (MGDA)
In this section, we briefly review the Multiple Gradient Descent Algorithm (MGDA)  [19], which serves as a fundamental tool of our method. MGDA extends the single objective steepest gradient descent algorithm for solving multi-objective optimization problems. It updates the parameters as follows: ${\bm{\theta}}_{k+1}={\bm{\theta}}_{k}+\eta{\bm{d}}$, where $\eta$ is the learning rate and ${\bm{d}}$ is the MGDA search direction. The key idea of MGDA is to find a valid gradient direction to decrease all objectives ${\bm{L}}_{i}({\bm{\theta}})$ at the same time. One way for computing ${\bm{d}}$ is as follows [20]:
@@
s.t. $\;\nabla{{\bm{L}}_{i}}^{T}{\bm{v}}\leq\alpha,\quad i\in[m].$
@@
If the optimal $\alpha^{*}<0$, then ${\bm{d}}$ is a direction that can reduce all the objectives, and each objective ${\bm{L}}_{i}$ can be decreased by about $\eta\alpha^{*}$. When $\alpha^{*}=0$, no direction can optimize all the objectives simultaneously, and thus ${\bm{\theta}}_{k}$ is Pareto stationary [19].
@@ III-A Preference Functions
III-A1 Exact preference function
There are several functions available to align a Pareto objective vector ${\bm{L}}({\bm{\theta}})$ with a preference vector ${\bm{\lambda}}$. One commonly used function is the Euclidean distance, which is initially introduced as part of the penalty boundary indicator (PBI) in Zhang and Li’s MOEA/D work [26]. The preference function based on Euclidean distance is defined as:
@@ III-A1 Exact preference function
This function indicates the exactness of a solution when $h({\bm{\theta}})$ is equal to zero.
III-A2 Region of Interest (ROI) Preference Function
The region of interest (ROI) is a specific region defined by the user. In addition to the regular ROI discussed in Section II-B, we also consider the ROI as a spherical shape. The preference function for the ROI constraint is defined as follows:
@@ III-B The Predict-and-Correct Framework
The objective of the preference-based multi-objective optimization (MOO) problem is to identify a Pareto optimal solution that fulfills the preference constraint $h({\bm{\theta}})=0$. Drawing inspiration from Gebken et al.’s work [27], we propose a predict-and-correct algorithmic framework. In this framework, let ${\bm{\theta}}_{k}$ represent the solution at iteration $k$, with the subsequent solution ${\bm{\theta}}_{k+1}$ determined as follows:
  • When $h({\bm{\theta}}_{k})=0$, we do the prediction step, which reduces all objective values ${\bm{L}}_{i}$ without hurting the preference constraint $h({\bm{\theta}}_{k})$:
    $\begin{split}({\bm{d}},\alpha^{*})=&\operatorname*{arg\,min}_{({\bm{v}},\alpha)\in\mathbb{R}^{m+1}}\alpha+\frac{1}{2}\left\|{\bm{v}}\right\|^{2},\\ \text{s.t.}\quad&\;\begin{cases}\nabla{{\bm{L}}_{i}}^{T}{\bm{v}}\leq\alpha,\quad i\in[m],\\ \nabla h^{T}{\bm{v}}=0.\end{cases}\end{split}$ (8)
  • Otherwise, we do the correction step. The major requirement is to reduce $h({\bm{\theta}}_{k})$ while the objective values ${\bm{L}}_{i}$ should not be deteriorated too much. Specifically, we compute the direction $\bm{d}$ as follows:
    $\begin{split}&({\bm{d}},\alpha^{*})=\operatorname*{arg\,min}_{({\bm{v}}\in\mathbb{R}^{n},\alpha\in\mathbb{R})}\alpha\\ &\text{s.t.}\quad\left\{\begin{aligned} &\nabla{\bm{L}}_{i}^{T}\bm{v}\leq\alpha,\quad i\in[m]\\ &\nabla h^{T}\bm{v}\leq-\sigma\left\|\nabla h\right\|\cdot\left\|\bm{v}\right\|,\\ &0<\left\|\bm{v}\right\|\leq 1.\\ \end{aligned}\right.\end{split}$ (9)
@@
We would like to offer some remarks regarding the conceptual algorithm we just introduced. During the prediction phase, there are two critical constraints to consider. The first one ensures that each objective, denoted as ${\bm{L}}_{i}$, can be reduced by a minimum of $\alpha$. The second constraint is designed to guide the search within the tangent space, ensuring that the resulting $h({\bm{\theta}}_{k})$ value remains minimal. It’s important to note that the Problem (8) is essentially a variation of Problem (4) and (5) with an additional quality constraint $\nabla h^{T}{\bm{v}}=0$. Therefore, the techniques used for analyzing and solving these equations can also be effectively applied to Problem (8).
In the correction phase, the first constraint outlined in Problem (9) aligns with the well-established Armijo rule (Armijo, 1966) [28]. This constraint is designed to ensure a sufficient reduction in the constraint function $h({\bm{\theta}})$. Such a reduction is pivotal for steering $h({\bm{\theta}})$ towards a stationary solution, a concept validated by Bertsekas in 2016 [29]. The second constraint regulates the increment of each ${\bm{L}}_{i}$ during the search process. The third constraint is critical in establishing the search boundary, ensuring that the vector $\bm{v}$ remains non-zero. When the correction step is exactly a retraction map as defined in  [30] onto the manifold $h({\bm{\theta}})=0$ at ${\bm{\theta}}_{k}$, our method is a special case of the constrained multiple gradient method introduced in  [27] and converges to a Pareto stationary solution satisfying the exact constraint.

### 16 Feb 2024 v2

@@ Abstract
It is desirable in many multi-objective machine learning applications, such as multi-task learning with conflicting objectives and multi-objective reinforcement learning, to find a Pareto solution that can match a given preference of a decision maker. These problems are often large-scale with available gradient information but cannot be handled very well by the existing algorithms. To tackle this critical issue, this paper proposes a novel predict-and-correct framework for locating a Pareto solution that fits the preference of a decision maker. In the proposed framework, a constraint function is introduced in the search progress to align the solution with a user-specific preference, which can be optimized simultaneously with multiple objective functions. Experimental results show that our proposed method can efficiently find a particular Pareto solution under the demand of a decision maker for standard multiobjective benchmark, multi-task learning, and multi-objective reinforcement learning problems with more than thousands of decision variables.
@@ I Introduction
Recently, Multi-Objective Optimization (MOO) concepts and methodologies have received growing attention in the machine learning community since many machine learning applications can be naturally modeled as MOO problems  [1, 2, 3]. For example, a face recognition system has to balance prediction accuracy on different attributes of the face [4, 5]. A recommendation system needs to balance several metrics, such as accuracy, novelty, and diversity of products [6, 7]. Very often, no single best solution can optimize all these objectives simultaneously. Instead of a single optimal solution, there exists a set of Pareto-optimal solutions with different optimal trade-off ratios.
Modern machine learning systems usually use deep neural networks as the core model. MOO problems in these machine learning systems [2, 8, 9, 5, 4] often share the following characteristics:
@@
TABLE I: Comparison of PMGDA with other gradient-based MOO methods.
@@
 EPO [14] Exact Pareto solutions.
(1) EPO is only designed for minimization tasks.
(2) EPO can only deal with positive objectives.
(3) Solution precision of EPO is bad on a number of MOO problems.
 COSMOS [4] Approximate exact Pareto solution. Solution quality is heavily depended on its control parameter $\mu$ and this parameter is difficult to tune.
 PMTL [8] Pareto solutions restricted in sector Areas. Challenging to apply for objectives $m$>2 due to sector divisions in 3-D (or more) space is difficult.
 mTche [15] Exact Pareto solutions.
(1) mTche is inefficient for MO problems with more than two objectives ($m>2$).
(2) Final performance is bad on complex MO reinforcement learning problems.
@@
Pareto solutions under inequality/
equality constraints.
PMGDA is particular designed for
exact and ROI Pareto solutions.
PMGDA solves a LP/QP (Similar to EPO/PMTL), taking more time than mTche.
Due to these characteristics, gradient-based methods are particularly well-suited for solving MOO problems in machine learning. It is, however, very costly to find a set of distinct Pareto solutions (e.g., deep neural networks) to approximate the entire Pareto front accurately. Some efforts have been made to design multi-objective gradient algorithms to find one or a few Pareto solutions [2, 8, 16, 3, 17] with successful applications on multi-task learning, image classification, face recognition, and other scene understanding tasks [18, 2] based on the Multiple Gradient Descent Algorithm (MGDA) [19]. However, a significant limitation of most of these gradient algorithms [2, 8, 19, 20] is their lack of consideration for the decision maker’s preferences.
With a very limited number of solutions (only one or a few), it is of great importance to develop algorithms for finding Pareto solutions satisfying user’s specialized demands. Some initial studies have explored approaches for finding Pareto objective solutions that align with user preferences. For instance, Pareto Multi-Task Learning (PMTL) [8] identifies Pareto solutions within specific sector regions, and the COSMOS method [4] incorporates a penalty term to approximate a Pareto solution near a preference vector. The Exact Pareto Optimization (EPO) method from Mahapatra et al. [14] focuses on locating solutions along a preference vector, claiming to find the ‘exact’ solution with preference. However, these methods are still nascent and have their own disadvantages. For example, PMTL [8] struggles with efficiency for more than three objectives, and COSMOS [4] is limited to approximate solutions. EPO employs a complex and specific search mechanism to divide objectives into different sets with different treatments. This approach sometimes fails to find the preferred solutions  [^2] If the linear programming problem defined in EPO fails, EPO relies on optimizing a linear combination of objectives. and performs poorly with more than two objectives. A comprehensive comparison with previous methods is presented in Table I.
Different from previous works, this paper introduces a general Preference-based Multiple Gradient Descent Algorithm, termed PMGDA, to effectively address user-specific preferences within the MGDA framework [19]. The preference assignment in PMGDA is flexible and general, which includes but is not limited to finding the exact Pareto solution [14]. The proposed PMGDA has a predict-and-correct approach to optimize the preference and all objective functions simultaneously. This paper makes the following main contributions:
@@
    We propose a novel framework for identifying Pareto solutions that align with a wide range of general preference functions, where finding exact Pareto solutions is a special case. Our approach is the first to effectively find Pareto solutions that meet specific region of interest (ROI) [21, 22] requirements.
@@
    We rigorously test PMGDA on well-known synthetic challenges, deep multi-task learning, and multiobjective reinforcement learning problems. PMGDA surpasses previous approaches in terms of solution precision and computational efficiency for finding exact Pareto solutions, while it is the first approach to successfully find Pareto solutions that satisfy the ROI constraint.
@@ II-A Notations
In this paper, the following notations are used (refer to Table II): Vectors are denoted using bold letters (e.g. ${\bm{v}}$). Matrices are denoted by an uppercase bold letters (e.g., ${\bm{G}}$). The index set $\{1,\ldots,m\}$ is shorthanded by $[m]$. We use three symbols to compare the relationship between two vectors ${\bm{x}}$ and ${\bm{y}}$:
For two vectors, ${\bm{x}}$ and ${\bm{y}}$, ${\bm{x}}\prec{\bm{y}}$ if $x_{i}<y_{i}$ for all $i\in[m]$. ${\bm{x}}\preceq{\bm{y}}$ if $x_{i}\leq y_{i}$ for all $i\in[m]$. ${\bm{x}}\preceq_{\text{strict}}{\bm{y}}$ if there exists $j\in[m]$ such that $x_{j}<y_{j}$ and $x_{i}\leq y_{i}$ for all $i\in[m]\setminus j$. In this article, $x_{i}$ denotes the $i$-th entry of vector ${\bm{x}}$, and ${\bm{x}}^{(i)}$ denotes the $i$-th vector in a set of vectors.
TABLE II: Notations used in this paper.
@@
$\Delta^{m-1}$ The (m-1)-simplex, $\left\{{\bm{y}}|\sum_{i=1}^{m}y_{i}=1,y_{i}\geq 0\right\}$.
@@
$\hat{\bm{x}}$ $\hat{\bm{x}}=\frac{{\bm{x}}}{\left\|{\bm{x}}\right\|}$. I.e, $\hat{\bm{x}}$ is a unit vector.
$\bm{\mu},\nu$ Lagrange multipliers.
@@ II-B Basic MOO Definitions
A MOO problem can be expressed as the following formulation:
$\min_{{\bm{\theta}}\in\Theta\subset\mathbb{R}^{n}}{\bm{L}}({\bm{\theta}})=(L_{1}({\bm{\theta}}),L_{2}({\bm{\theta}}),\ldots,L_{m}({\bm{\theta}})),$ (1)
where ${\bm{\theta}}$ is a $n$-D decision variable, and $\Theta$ is its decision space. Very often, no single solution can optimize all the objectives at the same time. Instead, there exists a set of solutions, called Pareto solutions, which can represent different optimal trade-offs among the objectives. The following MOO definitions are used in this paper. We have the following definition:
@@ A MOO problem can be expressed as the following formulation:
A solution is Pareto optimal if no other solution belongs to $\Theta$ dominates it. The set of all the Pareto solutions is called the Pareto set (PS); its image in the objective space is called the Pareto front ($\mathcal{T}$), $\mathcal{T}=f\circ\mathtt{PS}$.
@@
For objectives $L_{1},\ldots,L_{m}$ under the constraint $h({\bm{\theta}})$, a solution ${\bm{\theta}}$ is deemed Pareto stationary if it satisfies the following condition: there are no vectors $\mu_{1},\ldots,\mu_{m}$, and $\nu$ (not all zero) for which
@@
(a) Exact Pareto solution.
(b) Pareto solutions with a region of interest (ROI).
Figure 1: Two particular cases of preference-based MOO. The left figure: the exact Pareto solution aligns with a given preference vector. The right figure: ROI Pareto solutions satisfy the ROI constraint.
A solution is Pareto stationary means that, this solution can not be locally improved, which is a necessary condition for Pareto solution. Since a Pareto set could contain more than one single solution, and the possible infinite Pareto solutions are not comparable (e.g., non-dominated) with each other, therefore, it is necessary to make a further decision among the Pareto solutions [24]. Preference-based MOO aims to search for a Pareto solution that satisfies a user-defined preference function $h({\bm{\theta}})=0$. A direct application is to search the so-called ‘exact’ Pareto solutions.
@@
Let ${\bm{\lambda}}=(\lambda_{1},\ldots,\lambda_{m})$ be a positive preference vector, i.e., $\sum_{i=1}^{m}\lambda_{i}=1$ and $\lambda_{i}>0$, $\forall i\in[m]$. A solution ${\bm{\theta}}$ is called an ‘exact’ Pareto solution [14] for the preference ${\bm{\lambda}}$ if this solution is Pareto optimal and satisfies:
$\frac{L_{1}({\bm{\theta}})}{\lambda_{1}}=\frac{L_{2}({\bm{\theta}})}{\lambda_{2}}=\ldots=\frac{L_{m}({\bm{\theta}})}{\lambda_{m}}.$ (3)
Another application of preference-based MOO is that, in many cases, particularly when dealing with large problem sizes, the user may only be interested in a specific region of interest (ROI) on the Pareto front [21]. This falls under the category of preference-based MOO [22]. Previous literature has defined the ROI using the reservation point $\overline{\bm{y}}$ and the aspiration point $\underline{\bm{y}}$. The ROI can be represented as $\mathtt{ROI}=\{\bm{y}|\underline{\bm{y}}\preceq\bm{y}\preceq\overline{\bm{y}}\}$ [25]. Illustrations of ‘exact’ solution and ROI are shown in Figure 1.
@@ II-C The Multiple Gradient Descent Algorithm (MGDA)
In this section, we briefly review MGDA  [19], which serves as a fundamental tool of our method. MGDA extends the single objective steepest gradient descent algorithm for solving MOO problems. It updates the parameters as follows: ${\bm{\theta}}_{k+1}={\bm{\theta}}_{k}+\eta{\bm{d}}$, where $\eta$ is the learning rate and ${\bm{d}}$ is the MGDA search direction. The key idea of MGDA is to find a valid gradient direction to decrease all objectives $L_{i}({\bm{\theta}})$ at the same time. One way for computing ${\bm{d}}$ is as follows [20]:
@@
s.t. $\;\nabla{L_{i}}^{\top}{\bm{v}}\leq\alpha,\quad i\in[m].$
@@
If the optimal $\alpha^{*}<0$, then ${\bm{d}}$ is a direction that can reduce all the objectives, and each objective $L_{i}$ can be decreased by about $\eta\alpha^{*}$. When $\alpha^{*}=0$, no direction can optimize all the objectives simultaneously, and thus ${\bm{\theta}}_{k}$ is Pareto stationary [19].
@@ III-A Preference Functions
III-A1 The exact preference function
There are several functions available to align a Pareto objective vector ${\bm{L}}({\bm{\theta}})$ with a preference vector ${\bm{\lambda}}$. One commonly used function is the Euclidean distance from a Pareto objective to this preference vector [26].
@@ III-A1 The exact preference function
This function indicates a solution is exact if and only if $h({\bm{\theta}})$ is equal to zero.
III-A2 The Region of Interest (ROI) preference function
ROI is a specific region defined by the decision maker. In addition to the regular ROI discussed in Section II-B, we also consider the ROI as a spherical shape. The preference function for the ROI constraint is defined as follows:
@@ III-B The Predict-and-Correct Framework
The objective of the preference-based MOO (MOO) problem is to identify a Pareto solution that fulfills the preference constraint $h({\bm{\theta}})=0$. Drawing inspiration from Gebken et al.’s work [27], we propose a predict-and-correct algorithmic framework. In this framework, let ${\bm{\theta}}_{k}$ represent the solution at iteration $k$, with the subsequent solution ${\bm{\theta}}_{k+1}$ determined as follows:
  • When $h({\bm{\theta}}_{k})=0$, we do the prediction step, which reduces all objective values $L_{i}$ without hurting the preference constraint $h({\bm{\theta}}_{k})$:
    $\begin{split}({\bm{d}},\alpha^{*})=&\operatorname*{arg\,min}_{({\bm{v}},\alpha)\in\mathbb{R}^{m+1}}\alpha+\frac{1}{2}\left\|{\bm{v}}\right\|^{2},\\ \text{s.t.}\quad&\;\begin{cases}\nabla{L_{i}}^{\top}{\bm{v}}\leq\alpha,\quad i\in[m],\\ \nabla h^{\top}{\bm{v}}=0.\end{cases}\end{split}$ (8)
  • Otherwise, we do the correction step. The major requirement is to reduce $h({\bm{\theta}}_{k})$ while the objective values $L_{i}$ should not be deteriorated too much. Specifically, we compute the direction $\bm{d}$ as follows:
    $\begin{split}&({\bm{d}},\alpha^{*})=\operatorname*{arg\,min}_{({\bm{v}}\in\mathbb{R}^{n},\alpha\in\mathbb{R})}\alpha\\ &\text{s.t.}\quad\left\{\begin{aligned} &\nabla L_{i}^{\top}{\bm{v}}\leq\alpha,\quad i\in[m]\\ &\nabla h^{\top}{\bm{v}}\leq-\sigma\left\|\nabla h\right\|\cdot\left\|{\bm{v}}\right\|,\\ &0<\left\|{\bm{v}}\right\|\leq 1.\\ \end{aligned}\right.\end{split}$ (9)
@@
We would like to offer some remarks regarding the conceptual algorithm we just introduced. During the prediction phase, there are two critical constraints to consider. The first one ensures that each objective, denoted as $L_{i}$, can be reduced by a minimum of $\alpha$. The second constraint is designed to guide the search within the tangent space, ensuring that the resulting $h({\bm{\theta}}_{k})$ value remains minimal. It’s important to note that the Problem (8) is essentially a variation of Problem (4) and (5) with an additional quality constraint $\nabla h^{\top}{\bm{v}}=0$. Therefore, the techniques used for analyzing and solving these equations can also be effectively applied to Problem (8).
In the correction phase, the first constraint outlined in Problem (9) aligns with the well-established Armijo rule (Armijo, 1966) [28]. This constraint is designed to ensure a sufficient reduction in the constraint function $h({\bm{\theta}})$. Such a reduction is pivotal for steering $h({\bm{\theta}})$ towards a stationary solution, a concept validated by Bertsekas in 2016 [29]. The second constraint regulates the increment of each $L_{i}$ during the search process. The third constraint is critical in establishing the search boundary, ensuring that the vector ${\bm{v}}$ remains non-zero. When the correction step is exactly a retraction map as defined in  [30] onto the manifold $h({\bm{\theta}})=0$ at ${\bm{\theta}}_{k}$, our method is a special case of the constrained multiple gradient method introduced in  [27] and converges to a Pareto stationary solution satisfying the exact constraint.